### v1

# Designing and Evaluating AI Margin Notes in Document Reader Software

Abstract.
AI capabilities for document reader software are usually presented in separate chat interfaces. We explore integrating AI into document comments, a concept we formalize as AI margin notes. Three design parameters characterize this approach: margin notes are integrated with the text while chat interfaces are not; selecting text for a margin note can be automated through AI or manual; and the generation of a margin note can involve AI to various degrees. Two experiments investigate integration and selection automation, with results showing participants prefer integrated AI margin notes and manual selection. A third experiment explores human and AI involvement through six alternative techniques. Techniques with less AI involvement resulted in more psychological ownership, but faster and less effortful designs are generally preferred. Surprisingly, the degree of AI involvement had no measurable effect on reading comprehension. Our work shows that AI margin notes are desirable and contributes implications for their design.
interaction techniques, controlled experiments, large language models, generative AI, note-taking
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; ; copyright: noneccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Interaction techniques
Figure 1. Different ways to leverage LLM capabilities in document reader software: (a) we propose AI margin notes that are integrated with the document text unlike (b) chat-based interfaces that are separated from the document text.
[…]
Side-by-side comparison of two interfaces for describing text about flowering plants. Left: AI margin note highlights text, providing an explanation beside the document. Right: Chat-based interface with a chatbot explaining the same concept. Both focus on floral structures with a modern, educational tone.
1. Introduction
Taking notes while reading documents is a common active reading strategy that can be done in two primary ways. First, people can write notes that are decoupled from the document text by writing on a separate piece of paper. Second, people can write notes that are integrated with the document text (Adler et al., 1998), for example, by writing in the margins (marginalia). Such ‘margin notes’ have been common practice for centuries, as they provide a space for readers to summarize, paraphrase, explain unfamiliar concepts, make connections to existing knowledge, and even express personal opinions (Jackson, 2001; Porter-O’Donnell, 2004). This can be especially beneficial when reading for educational purposes (Bold and Wagstaff, 2017), as it provides space for people to work through difficult sections and quickly re-access their thoughts afterwards within a shared context (Marshall, 1997; Adler et al., 1998). When reading digital documents, margin notes can be created using pen input devices (e.g., (Romat et al., 2019)), but a more common way is by leaving comments in systems like Google Docs, Microsoft Word, and Adobe Acrobat Reader (Wolfe and Neuwirth, 2001). This is usually done by selecting text within a document and writing in a text box positioned beside it.
More recently, people have been leveraging the capabilities of large language models (LLMs) to support reading documents, for example, to summarize, reword, and better understand unfamiliar concepts (Guidroz et al., 2025; Gu et al., 2024; Kreijkes et al., 2025). The generated output serves a purpose like margin notes, yet the capabilities of LLMs have not been integrated into commenting features of document reader software. Instead, most LLM-enhanced document readers require users to summarize and ask questions about the text in a separate, chat-like interface disconnected from the document. There is an opportunity to improve how LLMs are used in this context by integrating their capabilities directly into ‘margin note’ comments that are linked to the document text.
We explore the design of “AI margin notes” that leverage LLMs to enhance comments in document reader software through three controlled experiments. Every experiment required participants to read short, non-fiction documents while interacting with an LLM, primarily within comments, and answer reading comprehension questions two hours later. Each experiment focused on a different design parameter. First, we compared AI margin notes to traditional, chat-based prompting to better understand the effects of integration. Second, we compared manually selecting text that an AI margin note is associated with, to automatically selecting text through an LLM-powered assistant to better understand the effects of selection automation. Third, we explored different AI margin note techniques that leverage LLMs in different ways to understand the effects of human and AI involvement. Specifically, AI margin notes that generate summaries, fill-in-the-blank exercises, responses to specific prompts, practice short answer questions, and feedback on how written text can be improved.
Our results indicated that AI margin notes were preferred over chat interfaces and that selecting text for AI margin notes should be done manually. AI margin note techniques with varying levels of human and AI involvement were valued for different reasons. For example, techniques with more human involvement were typically associated with more psychological ownership, but techniques with more AI involvement were faster, less effortful, and generally preferred. Our work contributes the idea of AI margin notes and empirical results demonstrating that AI margin notes are a desirable feature for document reader software.
2. Background and Related Work
AI margin notes relate to existing literature in psychology about the benefits and challenges of note-taking while reading, and other techniques that have used LLMs to improve reading and note-taking. We focus specifically on relevant work focused on reading, rather than taking notes while attending a lecture or watching a video.
2.1. Note-Taking while Reading
There are many types of notes that readers can create, margin notes being one of them. Research in psychology suggests that the general activity of note-taking is beneficial for two main reasons (Kiewra, 1989). First, notes act as external storage for information, which can be re-read to reinforce memory (the storage function). Second, the act of creating notes can facilitate learning as it requires readers to pay more attention to the material to process and organize ideas (the encoding function).
The encoding function is especially beneficial when the reader processes the text at a deeper level (Kiewra, 1985; Craik and Lockhart, 1972), for example, by connecting it to prior knowledge and experiences. However, many people opt for shallower and less effective note-taking strategies. For example, Bretzing and Kulhavy’s analysis of students’ note-taking activities (Bretzing and Kulhavy, 1981) revealed that students tend to take verbatim notes that repeats text from documents. In another study (Bretzing and Kulhavy, 1979), they showed that students who took verbatim notes performed worse on a test than those who used deeper note-taking strategies by writing their own summaries or by paraphrasing text from the document.
Note-taking is a complex activity that requires significant cognitive effort as readers must coordinate and frequently switch between reading and writing activities. Writing takes more time than reading, and excessive delays between reading activities as a result of note-taking can hinder comprehension. Therefore, readers often experience significant mental and temporal demand, even when they are reading and taking notes without any time limits (Piolat et al., 2005). This can become especially tiring with longer (Kobayashi, 2005) or poorly-formatted documents (Olive and Barbier, 2017). Although researchers have argued that increased cognitive effort can benefit learning (Bjork, 1994b; Bjork and Bjork, 2011; Bjork, 1994a) and memory (Tyler et al., 1979), it does not always lead to improved learning outcomes, for example, if the task is too frustrating or if the learner lacks motivation (Grund et al., 2024).
2.2. Reading and Note-Taking with LLMs
Recent work has investigated how LLMs can improve user experience and comprehension while reading and taking notes. For example, text simplification techniques that turn complex documents into simplified versions (Guidroz et al., 2025) and make documents easier to skim (Gu et al., 2024) can improve reading comprehension and reduce workload. Users of systems like ChatGPT, Adobe Acrobat (Adobe, 2025), Google Notebook LM (Google, 2025), NoteGPT (NoteGPT, 2025), and ChatPDF (ChatPDF, 2023) can ask questions, summarize, and write notes about documents. However, few investigations examine the effect on factors like reading comprehension, reading duration, workload, and preferences.
Kreijkes et al. (Kreijkes et al., 2025) asked high school students to try different note-taking techniques to study for reading comprehension tests that took place three days later: note-taking independently and note-taking while having access to an LLM-powered chatbot to ask questions. Both of these note-taking techniques were compared to a baseline of asking the chatbot questions about the document, without any note-taking. Their results showed that both note-taking techniques led to better performance than just asking the chatbot questions about the document, suggesting that using LLMs in a more cognitively engaging way (i.e., with some note-taking) can improve reading comprehension. However, they did not compare the two note-taking conditions, making the effect of LLM use on note-taking unclear.
Although Kreijkes et al. found that just asking the LLM questions led to poorer test performance, it was preferred by participants as it was perceived to be more enjoyable and less effortful. Such findings have been reproduced in other studies focused on the effects of LLMs on learning more broadly. Systems like ChatGPT can allow for personalized learning experiences, which can improve academic performance, reduce mental effort, and motivate learners (Wang and Fan, 2025; Deng et al., 2025), however, learners may over-rely on these systems (Wang and Fan, 2025). A balance may be to encourage more cognitive engagement. For example, ChatPRCS (Wang et al., 2024) generates practice reading comprehension questions for students, which increased mental load but improved reading comprehension. Similarly, CoAsker (Liu et al., 2023) encourages students to generate practice questions with an LLM-powered assistant, which were displayed in a side panel much like notes beside the document. When compared to generating practice questions without assistance, receiving questions that were generated by the assistant led to higher reading comprehension scores.
To our knowledge, no prior work has thoroughly explored integrating LLMs into the commenting feature of document reader software. The closest prior work are prototype mockups in Melin-Higgins’ bachelor’s thesis (Melin-Higgins, 2024), in which users select document text to issue pre-determined prompts to a hypothetical LLM-powered assistant. Responses appear as ‘sticky notes’ beside the selected text, like margin notes, or they appear underneath the selected text by modifying the document structure. A very small 3-person study showed a preference for the margin note style. The study also only focused on usability and user preferences, and did not consider factors like reading comprehension. Furthermore, the prototypes were partially functioning Figma mockups, and none explored different levels of human and AI involvement.
Based on this research, AI margin notes could hinder or help readers. Creating them may limit deeper processing of the document text and lower comprehension. However, some AI margin note creation techniques may make note-taking less mentally demanding, so readers focus more on the underlying text while reading. Some techniques may even help readers take non-verbatim notes while improving motivation and cognitive engagement. Therefore, it is important to understand the effect of AI margin notes, and note-taking with LLMs more generally, on factors like test performance, workload, and user preferences. Our work contributes these important insights, which have been lacking in existing literature.
3. AI Margin Notes
We focus on three design parameters of AI margin notes: integration, selection automation, and the level of human and AI involvement.
3.1. Integration
The primary difference between AI margin notes and chat-based tools is integration. Specifically, AI margin notes enhance comments in document readers. Comments are anchored to specific text, making them integrated into the document, which contrasts with chat-based tools that are placed in a separate side panel with the content disconnected from the context (Figure 2). A separate interface for interacting with an AI assistant may suffice for general questions about the document or to receive overall summaries. However, readers also direct the AI assistant to specific parts of the text to contextualize their prompt (Kreijkes et al., 2025). Separating these responses from the text can be inefficient.
First, referring to specific parts of a document while prompting requires additional work (Masson et al., 2024). For example, if the user wants to simplify the language of a paragraph in a scientific document (e.g., (Guidroz et al., 2025)), they must first type a command like “simplify this,” select the paragraph from the document, then copy and paste it after the command. This copy and paste step could be avoided entirely if the user selected the paragraph and typed “simplify this” in an AI margin note. Or easier, by just selecting text and pressing a button that reifies the “simplify this” prompt (Masson et al., 2024; Beaudouin-Lafon and Mackay, 2000).
Second, shifting attention to a separate interface may distract from reading. Note-taking requires frequent switching between reading and writing (Adobe, 2025; Piolat et al., 2005), and researchers suggest that note-taking should “interrupt reading as little as possible” (Marshall, 1997). Having readers frequently switch between separate reading and writing interfaces may further increase cognitive load (Harrison et al., 1995). These switching costs could be mitigated by presenting the prompting interface alongside the specific text the user is reading.
Third, referring back to specific responses requires additional work. For example, suppose a user used a chat-based LLM interface to get an explanation for an unfamiliar concept in a document. The explanation may be remembered in the short term, but if the user revisits the document much later, they must scroll through a lengthy and poorly organized chat history to find it, which may be especially difficult for readers with a lower working memory capacity to do (Sanchez and Wiley, 2009). Linking responses directly to the relevant text avoids this issue. Some systems, like Adobe Acrobat Reader, support clicking on chat responses to highlight relevant parts of the document, however, this still requires scrolling through chat history. Alternatively, the user could re-issue the prompt, but this wastes time and resources. Of course, users could copy responses they wish to save and paste them in comments (Kreijkes et al., 2025), but this requires additional copy and paste steps that would not be needed with an AI margin note.
Figure 2. Integration: (a) a chat-based interface is not integrated since it is separated from the associated document text and (b) AI margin notes are integrated since they are associated to specific document text. In this and following two figures: green denotes human-written text, like a prompt, and blue represents AI-generated text, like a response. Yellow indicates text that the AI margin note is linked to.
[…]
Conceptual diagram comparing two layouts. (a) Separated: text on left, colored boxes on right. (b) Integrated: text and boxes combined, aligned side by side.
3.2. Selection Automation
Figure 3. Selection automation: (a) text can be automatically selected and multiple AI margin notes created at once and (b) the user can manually select text to create an AI margin note.
[…]
Conceptual diagram comparing two methods: (a) Automated summarization with a ”Summarize” button, arrow to a highlighted text box. (b) Manual with text selection, ”Add Summary” button, and similar output.
Figure 4. Levels of human and AI involvement: traditional margin notes consist of text that is 100% written by a human, but an AI margin note could consist of text that was 100% generated by AI, or could involve more human-generated text.
[…]
Conceptual diagram showing a shift from 100% AI with blue notes to 100% human with green notes, labeled ”AI margin notes with more human involvement.”
AI margin notes are associated with specific text, which defines the context for the prompt and a location to display the note. Specifying text for the note could be done automatically by the AI assistant or manually by the user (Figure 3). For example, consider summarization, a common task readers do in conventional margin notes (Jackson, 2001; Porter-O’Donnell, 2004) and with LLM systems (Melin-Higgins, 2024; Kreijkes et al., 2025). When using LLMs in document reader software, readers can generate a summary of the entire document automatically by pressing a single button (Adobe, 2025), or they can manually specify which parts of the document need to be summarized by copying and pasting text into the chat. When the reader just presses a button, the LLM decides which information is important and worth including in the summary, but when the reader copies text to summarize, they are deciding which parts of the document are most important. Prior work on text highlighting suggests that this decision process could improve learning outcomes and reading comprehension, but it could also increase mental effort (Yue et al., 2015; Joshi and Vogel, 2024).
Automating the selection of text to create AI margin notes may also impact psychological ownership, feelings of the learning experiences belonging to them (Pierce et al., 2001). Prior work shows that fostering psychological ownership, which can be achieved by giving learners more control over their learning (Pierce et al., 2001, 2003), can encourage more active learning (Griffith, 1990) and motivates learning (Soslau and Yost, 2007). Automatically generating AI margin notes reduces effort, but potentially at the risk of lowering comprehension and psychological ownership.
3.3. Human and AI Involvement
Currently, commenting in document reader software is intended for text that is 100% written by a human. Specifically, the user must formulate their own ideas and type into a text box. At the other extreme, pressing a “Summarize” button in a document reader (Adobe, 2025) produces text that is 100% written by the LLM and does not require any guidance from the human. However, there are techniques for producing text that require involvement from both the human and the LLM. For example, chat-based interfaces require the user to write a prompt so the LLM can produce relevant output text. Here, the user exerts control over what the output should be, and the LLM produces it (Joshi and Vogel, 2025b). These roles can be reversed too, for example, the LLM could ‘prompt’ the user to respond to practice reading comprehension questions (Wang et al., 2024; Liu et al., 2023).
Based on these existing variations of human and AI involvement when interacting with LLMs, we believe that AI margin notes could also vary in how much human and AI involvement they require to produce the final comment text (Figure 4). Techniques that require more human involvement likely require more cognitive engagement than those with more AI involvement, but this may be beneficial for reading comprehension (Bjork, 1994b; Bjork and Bjork, 2011; Bjork, 1994a; Tyler et al., 1979), may discourage readers from taking verbatim notes (Bretzing and Kulhavy, 1979), and may even improve feelings of psychological ownership (Joshi and Vogel, 2025b, a). However, frustrating cognitive engagement can discourage learners (Grund et al., 2024), which may be mitigated by increasing AI involvement.
4. Experimental Method
We conducted a series of experiments to explore the design parameters of AI margin notes. Specifically, we were interested in understanding how integration, selection automation, and human and AI involvement affect reading comprehension and duration, psychological ownership, task workload, and user preferences. All experiments were conducted using the Prolific crowdsourcing platform. [^1] https://www.prolific.com
4.1. Task
The primary experimental task was composed of two stages. First, a reading stage, where participants read non-fiction documents that were approximately 500 words each. For each document, they interacted with an LLM assistant using a different technique that represents a variation within the design parameter under evaluation. Second, a test stage, where participants completed reading comprehension tests about each document. Each test consisted of six multiple choice questions. The documents and questions were developed by Wallace et al. (Wallace et al., 2022) and were designed by a learning and reading specialist to be suitable for an eighth grade reading level. [^2] Note that we slightly modified the wording of some questions, as we noticed that some questions could be successfully answered by looking at the wording of other questions. At the time of writing this paper, Wallace et al.’s license agreement permits reuse, modification, public display, and redistribution for non-commercial research purposes. This is representative of documents targeted for the general public (Nielsen, 2015).
4.2. Apparatus
Our experimental system was a custom Node.js and React web application that implemented two types of interfaces, one for each stage: the reading interface, where participants read a document and interacted with an LLM using a specific technique (Figure 5); and the testing interface, where participants answered time-bounded, multiple-choice reading comprehension tests.
4.2.1. Reading Interface
The document was displayed on the left side of the screen. The specific design variation of the experimental condition was displayed beside it (details provided in each individual experiment). All design variations that involved an LLM used GPT 4.1 mini. The top of the screen contained a toolbar indicating how many comments or responses they still had to complete and a blue “Finish Reading” button to end the trial.
4.2.2. Testing Interface
The testing interface displayed six multiple choice questions at the centre of the screen. At the top of the screen was a toolbar that displayed how many questions the participant had answered, a countdown timer (displayed numerically and as a progress bar that shrank every second), and a blue “Finish Test” button that could be pressed to end the test.
Figure 5. Experimental reading interface: (a) toolbar containing instructions, (b) document to read, and (c) space for specific design variations to be displayed.
[…]
Interface with a header prompting users to create comments. On the left, text titled ”Flowering Plants” explains their structure. Right side has a blank space.
4.3. Procedure
Participants received a link to an experimental web application through Prolific. They had to use a desktop or laptop computer, which was strictly enforced through the web application. First, they entered basic demographic information and read instructions. Next, they completed the reading stage for all documents. There was no time limit for the reading stage, however, to focus on the effect of the different techniques rather than the number of responses, participants could only leave 3 AI margin notes (or receive 3 responses from the AI assistant, depending on the technique). After reading the document, they answered 10 questions about their experience and 1 question about their prior knowledge of the document’s content. They repeated this for the other documents, then answered 2 questions about their overall preferences. They also described other study aids they used for the test, such as taking a screenshot of the document or writing notes outside of the reading interface.
Two hours later, the participant was invited to return for the test stage. Given the lower reading difficulty of the documents, we increased the difficulty of the test by making it closed-book and restricted to 60 seconds.
4.4. Design
All experiments use within-subjects design since the documents were relatively short, and we wanted participants to compare their experiences with each technique. There is one independent variable, condition, which was randomly assigned. For a single condition, one of six documents from Wallace et al. (Wallace et al., 2022) was randomly-assigned. For increased internal validity, a single document was not restricted to a single condition and could be assigned to any condition.
Figure 6. Techniques tested in Experiment 1: (a) a chat-based interface and (b) AI margin notes.
[…]
Two interfaces: (a) a chat-based interface with a question about flower roles, and (b) an AI margin note with the same content displayed in a single comment box.
4.5. Measures
Reading comprehension and duration were calculated from logs, and all other metrics were calculated from subjective questionnaires that participants completed after the reading stages. All used a 0-100 interval range.
4.5.1. Reading Comprehension
This represents the number of questions that were correctly answered during the test stage (0-6 range). Analyzing Reading Comprehension gives insights into the effects of AI margin notes on overall learning, and whether techniques that encourage more human involvement improve learning outcomes.
4.5.2. Duration
This is the time taken in minutes to complete the reading stage. Some AI margin notes that require more manual effort or human involvement likely take longer to complete.
4.5.3. Psychological Ownership
This was measured by asking two quantitative questions about Personal Ownership and Responsibility (Caspi and Blau, 2011). The average of the two create a composite measure, as done in prior work (Joshi and Vogel, 2025a, b). The internal consistency reliability score, which describes how consistently different items on a questionnaire describe the same underlying concept, was high for all experiments (.80 $\leq\alpha\leq$ .90), suggesting that the composite measure was appropriate to use for data analysis. As described previously, this is an important measure since fostering psychological ownership can improve learner experiences.
4.5.4. Task Workload
We used factors from the NASA-TLX: Mental Demand, Physical Demand, Temporal Demand, Performance, Effort, and Frustration. Techniques that are more manual or that require more human involvement may involve more mental demand and effort, which could be beneficial for learning (Bjork, 1994b; Bjork and Bjork, 2011; Bjork, 1994a; Tyler et al., 1979), however, this may not happen if they are too frustrating (Grund et al., 2024).
4.5.5. Preferences
We asked about Frequency of Use, a question from the system usability scale (SUS) representing how frequently the participant would use the feature if made available to them. After trying all techniques, participants also gave each condition a Ranking, where 1 was the best, 2 was the second best, and so on. Ties were allowed. To establish an Overall Ranking of conditions, we use the Condorcet voting method (Young, 1988). An overall rank of 1 means that technique ‘defeats’ all others in pairwise comparisons. An overall rank of 2 means that technique ‘defeats’ all others except for the technique ranked 1, and so on. This is important to study, as users would likely want to keep using techniques that they like, which is especially useful if it also improves their learning outcomes.
5. Experiment 1: Integration
The goal of this experiment is to understand the effect of integration that is achieved with AI margin notes. Participants wrote prompts using a traditional, chat-based interface (chat), or an AI margin note (note).
5.1. Participants
We recruited 29 participants through Prolific. Participants were restricted to the United States and Canada, and those who had completed 2,500 previous tasks on the platform with a 99-100% approval rating. Three participants (10%) were removed for using other study aids prior to the test, leaving 26 valid responses (Table A.1). All self-reported proficiency in reading in English. Participants received \$10 in total, and the top 25% scoring participants received an additional \$5 bonus to incentivize them to do well on the test. The experiment took roughly 25 minutes in total.
5.2. Apparatus
The reading interface had two variations. For the chat-based prompting technique, a chat interface was displayed to the right of the document. Here, participants could type prompts in a text box and press a blue “Submit” button. Their prompts appeared like a chat message on the right in a grey bubble, and LLM responses were displayed on the left in a light blue bubble (Figure 6a).
To make more direct comparisons between the two variations, we had to create an AI margin note technique where participants type prompts and receive responses from an LLM. To create an AI margin note, the participant selected text in the document using their cursor, which caused a blue “Comment” button to appear. Clicking this created a text box to appear that was anchored to part of the document text and resembled a Google Docs comment (Figure 6b). The participant could type a prompt and press a blue “Confirm” button. After a few seconds, the LLM provided a response. The participant’s prompt appeared at the top of the comment in grey, italicized text with a dashed border around it, and the response appeared below in black text, prepended with a black ‘sparkle’ icon. This styling clearly distinguishes human-written text (no icon) from AI-written text (with a ‘sparkle’ icon) and main comment text (black text) from additional context that helped generate the comment (grey, italicized text with a dashed border). Participants could edit their prompt to receive new output and they could delete their comments. Clicking on individual comments highlighted the corresponding text in the document.
5.3. Results
We use Wilcoxon signed-rank tests to investigate the effects of condition on the various measures. To streamline the presentation of results, details of these statistical tests are shown in Table A.2.
Overall, there were no significant effects of condition on Reading Comprehension, Duration, Psychological Ownership, any of the workload-related factors, and Frequency of Use.
5.3.1. Preferences
Although there were no differences for the aforementioned measures, most participants preferred prompting with AI margin notes. From the Condorcet voting method, we observed that note received an Overall Ranking of 1, and chat received an Overall Ranking of 2. Notably, majority of participants (17, 65%) assigned note a Ranking of 1, and chat a Ranking of 2 (16, 62%).
To better understand why participants preferred note over chat, we examined participants’ free-form responses. Several (12, 46%) mentioned how note was easier and more intuitive than chat. Participants described how they “liked [not] having to move to a separate chat”(P21), how “selecting text [was] more natural when one has a question related to that part of the text”(P14), and how “having the comment embedded right into the document [was] handy and easy to reference along with the actual text”(P5).
Figure 7. Techniques tested in Experiment 2: (a) automatically selecting text to generate all AI margin notes at once and (b) manually selecting text and generating AI margin notes one-by-one.
[…]
Two processes for generating AI margin notes. In (a), automatic text selection generates notes after a button press. In (b), manual text selection generates notes from highlighted segments. Both methods produce concise notes summarizing flower structures and functions.
Overall, our results suggested that the integration of AI margin notes is highly desirable and preferable over chat-based interfaces as it prevented switching between interfaces, it was easier for readers to indicate text they had questions about, and it was to reference individual comments later. In this experiment, associating text to the AI margin note was done manually by the participant. However, selecting text could also be done automatically by the LLM, which we explore further in the following experiment.
6. Experiment 2: Selection Automation
The goal of this experiment is to understand the effect of selection automation when creating AI margin notes. Participants either manually selected what text should be summarized (manual), or pressed a button to let the AI assistant place three summary AI margin notes in the document (automatic).
6.1. Participants
Using the same inclusion criteria as Experiment 1, we recruited 32 new participants on Prolific. Two participants (6%) were excluded for not attempting to answer any reading comprehension questions, or for using other study aids, leaving 30 valid responses (Table A.3). All self-reported English reading proficiency. Participants received \$10 as remuneration, with the top 25% receiving an additional \$5 bonus. The experiment took approximately 25 minutes in total.
6.2. Apparatus
Both variations focused specifically on producing AI margin notes that summarized text as this is a representative task that is currently done manually, by the user copying and pasting specific parts of a document into a chat-based interface, or automatically, by pressing a button. For the automatic selection technique, the top toolbar displayed a blue “Generate Comments” button in the top left corner (Figure 7a). Pressing this caused three AI margin note comments that were linked to the document text to appear to the right of the document. Each summarized specific selections from the document, and was displayed in black text prepended with a black ‘sparkle’ icon. As before, participants could click individual comments to highlight the corresponding text, but they could not delete or edit individual comments. Instead, they could press the “Generate Comment” button again to regenerate all comments.
The manual selection technique worked like the previous experiment (Figure 7b): the participant selected text from the document and pressed a blue “Comment” button to generate a summary comment (appearing after a few seconds). The styling of the comment was the same, except that participants could also delete or regenerate individual comments. The participant had to manually create three summary comments.
Figure 8. Experiment 2 results: (a) Duration, (b) Psychological Ownership, (c) Performance, (d) Effort, and (e) Frequency of Use.
[…]
Boxplot comparison of Manual and Automatic approaches across five measures: Duration, Psychological Ownership, Performance, Effort, and Frequency of Use. Blue data points overlay black boxplots, indicating data distribution and variance.
6.3. Results
As before, we use Wilcoxon signed-rank tests where applicable. Details of statistical tests are shown in Table A.4.
We did not observe any significant effect of condition on Reading Comprehension, so we focus our results on other metrics.
6.3.1. Duration
Participants were 39 seconds slower when they manually selected text (Figure 8a). There was a significant effect of condition on Duration, revealing that manual ($\textsc{mdn}\!=\!3.43$, $\textsc{iqr}\!=\!2.15$) was slower than automatic ($\textsc{mdn}\!=\!2.78$, $\textsc{iqr}\!=\!2.14$).
6.3.2. Psychological Ownership
Though slower, participants generally felt more psychological ownership when they manually selected text (Figure 8b). A significant effect of condition on Psychological Ownership revealed that participants felt more Psychological Ownership with manual ($\textsc{mdn}\!=\!50.25$, $\textsc{iqr}\!=\!46.62$) than automatic ($\textsc{mdn}\!=\!6.75$, $\textsc{iqr}\!=\!10.88$). This was supported by free-form responses like “[manually selecting text] requires me to focus more on the task, [which] gives me more responsibility”(P20).
6.3.3. Task Workload
Participants felt like they performed better when selecting text manually, despite it requiring more effort. Specifically, there was a significant effect of condition on Performance (Figure 8c), suggesting that participants felt like they were better at creating AI margin notes with manual ($\textsc{mdn}\!=\!88$, $\textsc{iqr}\!=\!24.25$) than automatic ($\textsc{mdn}\!=\!80.5$, $\textsc{iqr}\!=\!49.5$).
A significant effect of condition on Effort (Figure 8d) suggested that participants felt more Effort for manual ($\textsc{mdn}\!=\!35$, $\textsc{iqr}\!=\!41.25$) than automatic ($\textsc{mdn}\!=\!14$, $\textsc{iqr}\!=\!38.75$). This was supported by free-form responses like “in being able to highlight on my own, I also feel as if I did some of the work and not just nothing”(P25). We did not observe any differences between manual and automatic for Mental Demand, Physical Demand, Temporal Demand, and Frustration.
6.3.4. Preferences
Participants generally seemed to prefer generating AI margin notes manually. There was a significant effect of condition on Frequency of Use (Figure 8e), suggesting a preference to use manual ($\textsc{mdn}\!=\!79.5$, $\textsc{iqr}\!=\!47.5$) more frequently than automatic ($\textsc{mdn}\!=\!55$, $\textsc{iqr}\!=\!68.75$). For Overall Ranking, we observed that participants tended to prefer manual, which was ranked first. Specifically, 23 participants (77%) gave manual a Ranking of 1, and 21 (70%) gave automatic a Ranking of 2.
Majority (16, 53%) appreciated manual for the increased control it enabled, for example: “I felt a little more control [when] I was able to pick what I wanted to be summarized. [When] I just pressed one button, it summarized everything, but not necessarily the areas I wanted it to”(P22). A few (4, 13%) mentioned how manual provided more incentive to read the document, for example: “generating all summaries at once will encourage users not to actually read the document for themselves. Generating individual summaries after I selected text meant I [had] to actually read the document”(P10).
Overall, our results suggested that selecting text for an AI margin note is slower and requires more effort when done manually. However, this may be a worthwhile trade-off, as manually selecting text was associated with higher feelings of psychological ownership, better perceived performance, and was preferred by participants. In this experiment, the resulting AI margin note was always a summary, however, there are many other types of AI margin notes that can be created that involve the human and AI in different ways.
7. AI Margin Note Techniques
Figure 9. Techniques tested in Experiment 3: (a) receive a summary (100% AI-written text), (b) complete a fill-in-the-blank exercise, (c) write a prompt, (d) answer a practice question, (e) receive feedback on written text, and (f) write a comment without AI (100% human-written text).
[…]
A series of six comments about the parts of a flower. Each shows different activities: summary, fill-in-the-blank, prompt writing, answering a practice question, feedback on text, and writing without AI. The comments focus on identifying and explaining the parts and functions of flowers: sepals, petals, male parts, and female parts.
To fully explore the design parameter of human and AI involvement, we implemented six techniques that fall along a continuum, ranging from AI margin notes whose text was generated entirely by the LLM, to those whose text was written entirely by the user (Figure 9). All are created by selecting text manually and pressing a blue “Comment” button.
7.1. Receive Summary
The technique with the least human involvement is receiving a generated summary where 100% of the text is written by AI (Figure 9a). This is identical to the manual text selection technique presented in Experiment 2.
7.2. Complete Fill-in-the-Blank Exercise
A technique with slightly more human involvement is receiving a generated summary, but with some keywords omitted to resemble a fill-in-the-blank question (Figure 9b). Like the previous technique, the LLM produces a summary for the text selected by the user, shown in black text and prepended by a black ‘sparkle’ icon. Each summary contains one or two ‘blanks,’ displayed using dropdowns containing three options. The user must select the correct answer from the provided options. Underneath, the comment displays feedback to notify the user whether their responses are correct (shown in green with a check mark) or incorrect (shown in red with an ‘X’). The user can delete the comment or regenerate it.
7.3. Write Prompt
A technique with some human and some AI involvement is writing a custom prompt for the LLM to respond to (Figure 9c). Writing a prompt allows the user to exert more control over the output, for example, to achieve goals that go beyond summarization such as learning additional information about the text or simplifying it. This is identical to the AI margin note technique described in Experiment 1.
7.4. Answer Practice Question
Another technique with some involvement from both the human and the AI is answering a practice reading comprehension question that was generated by the LLM, which has been shown to help with learning (Wang et al., 2024; Liu et al., 2023). The LLM produces a practice short-answer question based on the selected text (Figure 9d), prepended with a grey ‘sparkle’ icon and shown using grey italicized text with a dashed border. The user types in their answer in a text box underneath and presses a blue “Confirm” button to submit their response. The submitted response is then shown in black text. After a few seconds, the user receives feedback from the LLM about their answer: correct responses are shown in green and with a check mark underneath the response, and incorrect responses are shown in red with an ‘X.’ The user can edit their responses to correct mistakes (and will receive new feedback accordingly), or can delete their comment.
7.5. Receive Feedback on Written Text
A technique that has a lot of human involvement is requiring users to write their own comment text, with LLM-generated feedback and suggestions on how their comment can be improved (Figure 9e). This can be beneficial, as many learners struggle to take effective notes (Bretzing and Kulhavy, 1979). The user can type within the provided text box and press “Confirm.” The comment text is shown as black text. After a few seconds, the AI assistant provides feedback on the text, prepended with a grey ‘sparkle’ icon and shown as grey, italicized text with a dashed border. A green check mark is shown if their comment text is good, and a red ‘X’ is shown if it is not. The user can edit their comment to receive updated feedback and can delete it.
7.6. Write Text without AI
The technique with the most human involvement is requiring the user to write their own comment, without any AI assistance (Figure 9f). This baseline resembles the behaviour of existing commenting systems, where 100% of the text is written by the user. The user can type within the provided text box and press “Confirm.” The comment is shown as black text. The user can edit or delete the comment.
Figure 10. Experiment 3 results: (a) Duration, (b) Psychological Ownership, (c) Mental Demand, (d) Performance, and (e) Effort.
[…]
Boxplot comparison of None, Feedback, Question, Prompt, Blank, and Summary across five measures: Duration, Psychological Ownership, Mental Demand, Performance, and Effort. Coloured data points overlay black boxplots, indicating data distribution and variance.
8. Experiment 3: Human and AI Involvement
The goal of this experiment is to investigate the effect of human and AI involvement. Participants read six documents, one for each AI margin note technique described above: receiving a summary (summary), completing a fill-in-the-blank exercise (blank), writing a prompt (prompt), answering a practice question (question), receiving feedback on written text (feedback), and writing without any AI assistance (none). The experiment was longer, approximately 90 minutes.
8.1. Participants
We recruited 36 new participants on Prolific, using the same inclusion criteria as the previous experiments. Four (11%) were excluded for experiencing technical issues, leaving 32 valid responses (Table A.5). All but one were proficient at reading in English. Participants received \$25 total, with an additional \$5 bonus if their total reading comprehension score was within the top 25%.
8.2. Results
Where applicable, we use Friedman omnibus tests and Wilcoxon signed-rank post hoc tests, with Holm’s corrections for multiple comparisons. Statistical test details are shown in Table A.6.
8.2.1. Reading Comprehension
Contrary to what prior work suggests, techniques with more human involvement were not associated with higher reading comprehension scores. Specifically, the differences between summary ($\textsc{mdn}\!=\!4$, $\textsc{iqr}\!=\!2.25$) and all other conditions (all $\textsc{mdn}\!=\!5$, $\textsc{iqr}\!=\!2$) are marginal ($p=.05$), and not statistically significant.
8.2.2. Duration
Participants were faster when they used techniques that did not require as much human involvement (Figure 10a). A significant effect of condition on Duration revealed that summary ($\textsc{mdn}\!=\!3.96$, $\textsc{iqr}\!=\!2.78$) and blank ($\textsc{mdn}\!=\!4.92$, $\textsc{iqr}\!=\!2.80$) were faster than all other techniques.
8.2.3. Psychological Ownership
Generally, participants felt more psychological ownership for techniques that required more human involvement (Figure 10b). A significant effect of condition on Psychological Ownership and post hoc tests revealed that none ($\textsc{mdn}\!=\!100$, $\textsc{iqr}\!=\!4.38$) and feedback ($\textsc{mdn}\!=\!97$, $\textsc{iqr}\!=\!9.34$) were associated with the highest Psychological Ownership, and summary ($\textsc{mdn}\!=\!23.75$, $\textsc{iqr}\!=\!60.88$) with the lowest.
8.2.4. Task Workload
Generally, techniques that required more human involvement were more mentally demanding, associated with poorer perceived performance, and more effortful. For Mental Demand, there was a significant effect of condition (Figure 10c) and post hoc tests showed that none ($\textsc{mdn}\!=\!61$, $\textsc{iqr}\!=\!47$) and feedback ($\textsc{mdn}\!=\!70.5$, $\textsc{iqr}\!=\!31$) were associated with higher scores than prompt ($\textsc{mdn}\!=\!40.5$, $\textsc{iqr}\!=\!52.25$), blank ($\textsc{mdn}\!=\!20.5$, $\textsc{iqr}\!=\!30.75$), and summary ($\textsc{mdn}\!=\!11.5$, $\textsc{iqr}\!=\!22.25$).
There was also a significant effect of condition on Performance (Figure 10d), with post hoc tests revealing that participants generally felt like they did better at the task with blank ($\textsc{mdn}\!=\!100$, $\textsc{iqr}\!=\!8.5$) than none ($\textsc{mdn}\!=\!90.5$, $\textsc{iqr}\!=\!23.5$), feedback ($\textsc{mdn}\!=\!95$, $\textsc{iqr}\!=\!15.25$), and summary ($\textsc{mdn}\!=\!100$, $\textsc{iqr}\!=\!20.75$).
For Effort, a significant effect of condition (Figure 10e) and post hoc tests showed that none ($\textsc{mdn}\!=\!62.5$, $\textsc{iqr}\!=\!42.75$) and feedback ($\textsc{mdn}\!=\!68$, $\textsc{iqr}\!=\!32.5$) had the highest scores, and summary ($\textsc{mdn}\!=\!10$, $\textsc{iqr}\!=\!36.75$) had the lowest. We did not observe significant differences between the different techniques for Physical Demand, Temporal Demand, and Frustration.
8.2.5. Preferences
There were no differences in Frequency of Use, with all medians being 50 or greater, suggesting that participants would like to use all techniques. For Overall Ranking, the Condorcet voting method revealed that overall, blank was ranked first, followed by question, summary, prompt, feedback, and none. This Overall Ranking suggests that participants generally preferred techniques with more AI involvement (blank, question, summary, and prompt) over those with more human involvement (feedback and none), for example: “the more involved AI was, the more I appreciated the help”(P15).
Considering the Ranking scores, 21 (66%) ranked blank within the top 3, and 20 (62%) placed question within the top 3. These two techniques were generally valued as they “had the greatest pedagogical heft”(P7) and “actively engage memory and reinforce understanding”(P4). Nineteen (59%) placed summary within the top 3, however, eight (25%) gave it a rank of 6, suggesting that opinions were more divided for summary. Those that ranked summary highly generally valued it for “[making] work so much easier”(P9), however, some noted how they “want to do things [and] feel accomplished”(P30), goals that were not as well-supported with summary. Fifteen (47%) placed prompt within the top 3 as it was “simple and easy”(P23) and “encouraged curiosity”(P4). In contrast, seventeen (53%) ranked feedback within the bottom 3, and 19 (59%) ranked none within the bottom 3. Both techniques required “too much effort”(P20), which could be “draining and mind-absorbing”(P25).
9. Discussion
Overall, our results show that participants preferred AI margin notes over traditional, chat-based interactions. When AI margin notes were created through manual text selections, participants took longer to read the document and this required more effort, but they felt more psychological ownership and felt like they did a better job as a result. These feelings may have contributed to the strong preferences they had for manually-created AI margin notes. In short, our results suggest that AI margin notes were valued for their integration, and are perceived to be especially valuable with manual text selection.
Regarding human and AI involvement, our results suggest that different techniques have their own strengths and weaknesses. For example, techniques that required more human involvement, like receiving feedback about written text, required more mental demand and effort, but were associated with higher feelings of psychological ownership than those with low human involvement, like receiving summaries. However, participants preferred techniques with more AI involvement. One promising technique was receiving generated summaries with fill-in-the-blank exercises, the most preferred technique. It was perceived as not too mentally demanding or effortful, still associated with moderate feelings of psychological ownership, and participants felt like they performed well when creating them.
The implications for design are to adopt the integrated AI margin note approach within document reader software, to support manual text selection, and to provide options of AI margin note technique that vary in human and AI involvement, depending on the user’s preferences and goals.
9.1. Reading Comprehension
Given the literature that suggests that increased cognitive engagement can increase learning, we were surprised to see that more cognitively demanding AI margin note techniques were not associated with higher reading comprehension scores. To confirm this is not due to one specific experimental protocol, we conducted other pilot experiments with different participants and key variations. For example, we ran experiments with 24 hour gaps between the reading and testing stages to ensure that the documents were not simply remembered in short-term memory, but we did not observe any significant differences. Participants self-declared their knowledge of the topics discussed in each document, so we tried omitting participants with higher background knowledge, and found similar results. Another possibility is a ceiling effect because the range of possible scores is low (0-6). To rule this out, we conducted another pilot experiment using documents and questions from Guidroz et al. (Guidroz et al., 2025), [^3] We emailed the authors and received permission to reuse their materials. enabling more granular (0-12) scores. These documents were also more difficult (university graduate level, rather than an 8th grade level). Even with this increased score granularity and more difficult reading task, we found similar results, with no significant differences between techniques. Therefore, we believe that less cognitively engaging techniques with more AI involvement, like receiving generated summaries, may not be as detrimental to reading comprehension as one may assume.
9.2. Limitations
Our results, especially results related to reading comprehension, may not hold after extended long-term use. For example, LLMs may hallucinate and produce incorrect AI margin notes, and people may come to over-rely by studying incorrect notes (Wang and Fan, 2025). Techniques with low human involvement may make students ‘lazy’ over time and impact their ability to take notes in situations where LLMs are unavailable.
Our results may not also apply to different types of documents and longer documents. For example, prior work on text highlighting suggests that readers should be able to distinguish between important and unimportant text (Fowler and Barker, 1974), however, readers may struggle to identify the most important information within a long and complex document about an unfamiliar topic. In these cases, automatically selecting text within the document to create AI margin notes may be preferred.
Some aspects of our experimental design may be lacking in ecological validity. Notably, we chose to keep the number of AI margin notes and chat responses constant across techniques, and participants could only interact with one technique at a time. We made these decisions for increased control, however, such restrictions would not exist in a real system, where users would create as many AI margin notes and in whatever way they wish.
We focused on the specific context of creating AI margin notes as a study aid. However, people leave comments in document readers for many other reasons, for example, to leave editorial notes. Other AI margin note designs may be better suited to these goals.
10. Conclusion and Future Work
We propose and explore the design of “AI margin notes” that leverage the commenting feature of document reader software to provide LLM capabilities in a way that is more integrated into document text. Three experiments evaluated variations from different design parameters: integration, selection automation, and human and AI involvement, and overall, participants valued having integrated AI margin notes and creating them manually. AI margin note techniques that involved the human and AI to different degrees were valued for different reasons, suggesting that document reader software should provide multiple variations to support different user goals and preferences.
Future work should incorporate AI margin notes into existing software, explore other types of AI margin notes, and investigate the effects on factors like reading comprehension over the long term and with different types of documents. Our work adds more evidence that chat-based interfaces are not the only way of interacting with LLMs, and that increased integration with document text is beneficial, especially when they are created manually and the trade-offs of human and AI involvement are considered.
Appendix A Appendix
Table A.1. Demographics for Experiment 1.
Gender Age Education
Men 16 25-34 4 High School 2
Women 10 35-44 14 Some University (no credit) 5
45-54 3 Bachelor’s Degree 14
55-64 3 Professional Degree Beyond Bachelor’s 1
65-74 2 Master’s Degree 3
Doctorate Degree 1
Document Reader Frequency Commenting Frequency LLM Frequency LLM Summarization Frequency
Daily 6 Daily 1 Daily 10 Daily 3
Weekly 11 Weekly 6 Weekly 9 Weekly 9
Monthly 7 Monthly 3 Monthly 5 Monthly 3
Less than Monthly 2 Less than Monthly 5 Less than Monthly 5 Less than Monthly 5
Never 11 Never 1 Never 6
[…]
Demographic information: gender, age, education, experience using document readers, commenting experience, LLM experience, and LLM document summary experience. The demographics cover a diverse range.
Table A.2. Statistical test results for Experiment 1.
Measure $W$ $p$ RBCRBC
Reading Comprehension 80 .81 .06
Duration 148 .50 .16
Psychological Ownership 82.5 .40 .21
Mental Demand 149 .72 -.08
Physical Demand 76 .68 .11
Temporal Demand 74.5 .63 -.13
Performance 70.5 .78 .08
Effort 105.5 .32 .24
Frustration 57 .12 -.40
Frequency of Use 77.5 .07 .44
[…]
Details of statistical tests. Significant tests are denoted with asterisks, and more asterisks indicate smaller p-values.
Table A.3. Demographics for Experiment 2.
Gender Age Education
Men 18 25-34 4 High School 4
Women 12 35-44 15 Some University (no credit) 7
45-54 8 Bachelor’s Degree 9
55-64 1 Master’s Degree 10
65-74 1
75+ 1
Document Reader Frequency Commenting Frequency LLM Frequency LLM Summarization Frequency
Daily 4 Daily 1 Daily 8 Daily 2
Weekly 10 Weekly 3 Weekly 12 Weekly 8
Monthly 7 Monthly 4 Monthly 5 Monthly 4
Less than Monthly 7 Less than Monthly 3 Less than Monthly 5 Less than Monthly 11
Never 2 Never 18 Never 2 Never 5
[…]
Demographic information: gender, age, education, experience using document readers, commenting experience, LLM experience, and LLM document summary experience. The demographics cover a diverse range.
Table A.4. Statistical test results for Experiment 2.
Measure $W$ $p$ RBCRBC
Reading Comprehension 88.5 .20 .30
Duration 102 .006 .56 **
Psychological Ownership 9 ¡ .001 .95 ***
Mental Demand 158.5 .46 .16
Physical Demand 31.5 .10 .48
Temporal Demand 100 .25 .28
Performance 52.5 .02 .58 *
Effort 84.5 .004 .61 **
Frustration 77.5 .19 -.33
Frequency of Use 63.5 .004 .64 **
[…]
Details of statistical tests. Significant tests are denoted with asterisks, and more asterisks indicate smaller p-values.
Table A.5. Demographics for Experiment 3.
Gender Age Education
Men 13 18-24 1 High School 7
Women 19 25-34 2 Some University (no credit) 2
35-44 7 Technical Degree 5
45-54 13 Bachelor’s Degree 14
55-64 4 Master’s Degree 4
65-74 4
75+ 1
Document Reader Frequency Commenting Frequency LLM Frequency LLM Summarization Frequency
Daily 6 Daily 1 Daily 13 Daily 3
Weekly 14 Weekly 6 Weekly 11 Weekly 12
Monthly 8 Monthly 4 Monthly 5 Monthly 3
Less than Monthly 3 Less than Monthly 8 Less than Monthly 3 Less than Monthly 10
Never 1 Never 12 Never 4
[…]
Demographic information: gender, age, education, experience using document readers, commenting experience, LLM experience, and LLM document summary experience. The demographics cover a diverse range.
Table A.6. Statistical test results for Experiment 3.
Measure $Q$ $p$ $W$
Reading Comprehension 10.87 .05 .07
(a) Duration 64.64 ¡ .001 .40 ***
(b) Psychological Ownership 74.76 ¡ .001 .47 ***
(c) Mental Demand 57.83 ¡ .001 .36 ***
Physical Demand 20.92 ¡ .001 .13 *** post hocs were n.s.
Temporal Demand 4.45 .49 .03
(d) Performance 16.82 .005 .10 **
(e) Effort 55.98 ¡ .001 .35 ***
Frustration 5.39 .37 .03
Frequency of Use 9.19 .10 .06
(a) Duration (b) P. Ownership (c) Mental Demand (d) Performance (e) Effort
comparisons p-value p-value p-value p-value p-value
none feedback .02 * .24 .52 1 .91
none question .56 ¡ .001 *** .06 1 .004 **
none prompt .45 .003 ** .02 * 1 .01 *
none blank .003 ** ¡ .001 *** ¡ .001 *** .005 ** ¡ .001 ***
none summary ¡ .001 *** ¡ .001 *** ¡ .001 *** 1 ¡ .001 ***
feedback question .18 .002 ** .007 ** 1 .01 *
feedback prompt .04 * .003 ** .008 ** 1 .01 *
feedback blank ¡ .001 *** ¡ .001 *** ¡ .001 *** .03 * .001 **
feedback summary ¡ .001 *** ¡ .001 *** ¡ .001 *** 1 ¡ .001 ***
question prompt .19 .44 .52 1 .91
question blank .001 ** .004 ** .05 .08 .03 *
question summary ¡ .001 *** ¡ .001 *** ¡ .001 *** 1 .004 **
prompt blank .03 * .24 .15 .07 .11
prompt summary .002 ** ¡ .001 *** .02 * 1 .02 *
blank summary .14 .003 ** .15 .04 * .04 *
[…]
Details of statistical tests. Significant tests are denoted with asterisks, and more asterisks indicate smaller p-values.